# A GPU-based Compressible Combustion Solver for Applications Exhibiting Disparate Space and Time Scales

### v1

@@ Abstract
High-speed chemically active flows present significant computational challenges due to their disparate space and time scales, where stiff chemistry often dominates simulation time. While modern supercomputing scientific codes achieve exascale performance by leveraging graphics processing units (GPUs), existing GPU-based compressible combustion solvers face critical limitations in memory management, load balancing, and handling the highly localized nature of chemical reactions. To this end, we present a high-performance compressible reacting flow solver built on the AMReX framework and optimized for multi-GPU settings. Our approach addresses three fundamental GPU performance bottlenecks: memory access patterns through column-major storage optimization, computational workload variability via a novel bulk-sparse integration strategy for chemical kinetics, and multi-GPU load distribution for adaptive mesh refinement applications. The solver adapts existing matrix-based chemical kinetics formulations to multigrid contexts. Using representative combustion applications including hydrogen-air detonations and jet in supersonic crossflow configurations, we demonstrate $2-5\times$ performance improvements over initial GPU implementations with near-ideal weak scaling across 1961-96 NVIDIA H100 GPUs. Roofline analysis reveals substantial improvements in arithmetic intensity for both convection ($\sim 10\times$) and chemistry ($\sim 4\times$) routines, confirming efficient utilization of GPU memory bandwidth and computational resources.
@@
[…]
@@ 1 Introduction
Other studies have explored GPU usage or adaptive mesh refinement (AMR) techniques for accelerating CFD codes (Sharma et al., 2024b; Bernardini et al., 2021; Lai et al., 2019; Barwey and Raman, 2021; Bielawski et al., 2023; Romero et al., 2020; Witherden et al., 2015). AMR is an approach that helps address the disparate space and time scales of compressible reacting flow simulations on large, yet fine, meshes by resolving only spatio-temporal regions that need higher resolution. For instance, the AMReX framework (Zhang et al., 2019) — a structured-grid AMR coding framework — provides a hardware-agnostic solution for CFD code development, easing code portability to a diverse set of exascale architectures. AMReX-based solvers such as PeleC (Henry de Frahan et al., 2023) have demonstrated remarkable scalability. However, achieving optimal GPU performance demands algorithmic restructuring and hardware-specific tuning beyond CPU-GPU porting strategies provided by AMReX.
@@ 2.2 Graphics Processing Units and Roofline Models
Graphics processing units (GPUs) can significantly accelerate CFD calculations because many underlying algorithms can be cast into a single-instruction, multiple-data (SIMD) pattern, which GPUs are designed to exploit (Flynn, 1972). While CPUs excel at complex control flow and low-latency operations, GPUs contain hundreds of processing cores termed streaming multiprocessors (SMs) designed to execute identical operations on different data elements simultaneously. This hardware design makes them well suited for grid- and particle-based simulations, where the same numerical operations are applied independently across millions of computational elements with minimal interdependence. However, effectively utilizing GPUs requires addressing fundamental differences from CPU programming. A kernel is a function that runs many times in parallel on a GPU. The kernel executes work across many parallel workers, termed threads. The threads work in unison within groups called warps (32 threads per warp on NVIDIA hardware) that must execute identical instructions in lockstep. When threads within a warp take different execution paths – a condition known as thread divergence – performance degrades significantly. Additionally, GPUs achieve peak performance only when memory accesses follow coalesced patterns, where consecutive threads access consecutive memory locations (CUDA, ).
In modern GPUs, performance is predominantly limited by memory bandwidth rather than arithmetic throughput (Nickolls and Dally, 2010). For instance, the NVIDIA H100 GPU (NVIDIA H100 GPU, ) used in this study achieves peak double-precision performance of 34 TFLOPs/s but has global memory bandwidth of only 3.35 TB/s. This imbalance means that most computational kernels become memory-bound, where the costs of data movement dominate the costs of arithmetic operations. Common optimization strategies include minimizing CPU-GPU data transfers, utilizing on-chip shared memory for data reuse, ensuring coalesced memory access patterns, and maximizing arithmetic intensity—the ratio of floating-point operations to bytes transferred (Schroeder, 2011).
@@
Modern roofline analysis tools, such as NVIDIA Nsight Compute (NVIDIA Corporation, 2024), automate the collection of performance metrics necessary for roofline analysis. These tools measure kernel execution time, floating-point operations, and memory transactions to calculate both arithmetic intensity and sustained performance. The resulting roofline charts provide immediate visual feedback on performance bottlenecks and guide optimization efforts by clearly indicating whether kernel performance is limited by computational throughput, or other architectural constraints.
@@ 2.3 Overview of AMReX
This section provides a brief overview of the AMReX framework, focusing on the key concepts and features most relevant to our GPU-accelerated combustion solver. For comprehensive documentation of AMReX’s full capabilities, readers are referred to Zhang et al. (2019) and the official documentation (AMReX Development Team, 2025a).
@@
The AMReX framework organizes computational cells into rectangular non-overlapping grids at different AMR levels. The coarsest level – level 0 or the base level – contains the base coarse grids that span the entire physical domain, while higher levels contain dynamically generated finer grids that cover only regions meeting user-defined refinement criteria. To ensure efficient multigrid performance and optimal load balancing, AMReX enforces a blocking factor constraint, requiring all grid dimensions to be divisible by a specified power-of-two value (e.g., $2^{k}$ where $k\geq 0$). This constraint enables efficient grid refinement algorithms and facilitates regular memory access patterns, which are essential for achieving optimal GPU performance. From a programming standpoint, a Fortran Array Box (FAB) object efficiently stores the data and attributes for each grid, and FABs at a given AMR level are collectively stored in a multiFAB object (AMReX Development Team, 2025a). multiFAB provides a high-level interface for computational scientists to perform efficient parallel operations, load balancing, and communication. Unlike unstructured adaptive methods, this block-structured approach preserves regular data access patterns, which are essential for achieving optimal GPU performance.
@@
Load balancing across multiple GPUs is a key challenge in multigrid solvers. In CPU settings, users can calculate computational costs for each MultiFab and provide these to AMReX’s load balancing algorithms, which then redistribute the grids across the available MPI ranks to achieve balanced workloads AMReX Development Team (2025b). However, this strategy becomes limited on GPUs due to the typical one-rank-per-GPU deployment, which drastically reduces the number of redistribution targets compared to CPU runs that may use dozens of ranks per node. Additionally, GPU performance is nonlinear with respect to the number of grids and cells processed, making cost estimation more complex. Load balancing becomes particularly important in scientific codes where physical features – and thus computational costs – can vary dramatically across the domain.
@@ 3.1 Memory-Based Optimizations
AMReX already provides several GPU-based strategies for memory management and access. For instance, the framework manages memory through pre-allocated memory arenas that reuse contiguous memory chunks to eliminate allocation overhead (AMReX Development Team, 2025c). By default, amrex.the_arena_is_managed is set to 0, allocating device-only memory. Setting it to 1 enables managed memory accessible from both the CPU and GPU, facilitating easier development but potentially incurring performance penalties due to page faults. To address the penalties, we disabled managed memory and use explicit memory management via The_Async_Arena() for temporary allocations while leveraging AMReX’s arena system for memory pools. We also determined the maximum problem size for a single H100 based on the total device memory requirements for various variables. The variables include, but are not limited to, thermo-fluid variables, chemistry variables, temporary variables for flux computation, and variables for ghost cells.
@@ 3.2 Chemical Kinetics Solver
To reduce kernel launch costs in both bulk and sparse integration, we launch a single kernel across all cells from all grids. This design choice sacrifices some parallelization strategies over others. To guide this explanation, refer to Figure 2, which showcases two implementations of kinetics on GPUs. In a single-grid solver with contiguous memory layout, we could parallelize across both cells and species ($N_{c}\times N_{s}$ threads), enabling the use of GPU-optimized libraries like cuBLAS (NVIDIA Corporation, 2025) for matrix operations, as seen in Figure 2. This approach was utilized by Barwey and Raman (2021) to achieve superior performance when working with (i) a single medium to large grid (see C) or (ii) a low grid count in a multi-grid setting (see E). This multi-kernel strategy also eliminates warp divergence by checking whether all cells are complete before continuing the while loop iteration.
@@ 4.1 Setup of the Hardware, Software, and Performance Metrics
We obtain all results reported in Section 5 by running the solver on a cluster with NVIDIA H100 GPUs (NVIDIA H100 GPU, ). The GPUs are peer-to-peer connected using 18 4th-generation NVLINKs, with a total peak bandwidth of 900 GB/s, supporting simultaneous bidirectional communication. Empirical bandwidth tests using sample CUDA scripts (Corporation, 2024) measured net unidirectional data transfer rates of $\approx 478$ GB/s between any two GPUs. GPUs are connected to Intel(R) Xeon(R) Platinum 8468 CPUs using PCIe 5th Gen communication technologies, with peak bandwidths of 64 GB/s. The simulations were executed with each CPU process bound to a single GPU, as required by the AMReX GPU implementation. Therefore, the number of CPUs used for a simulation equals the number of GPUs.
We compiled two versions of the executable for a given set of inputs. The first version is compiled with level 3 optimizations using the GNU suite of compilers. The second version is a suboptimal compilation with debugging enabled, allowing for the collection of wall timing for various code regions. We also used the debugging version to collect detailed GPU kernel metrics using NVIDIA Nsight Compute (NVIDIA Corporation, 2024) to derive quantities for the roofline model. These metrics are detailed in Yang et al. (2020).
@@ 4.3 Description of The Combustion Applications
Jet in Supersonic Crossflow (JISCF): This application simulates a sonic Ethylene jet injected into a high-temperature Mach 2.48 crossflow, representing conditions relevant to high-speed propulsion systems. The interested reader may find details of the case in Sharma et al. (2024c). The configuration features complex physics, including compressible turbulence, shock-turbulence interactions, fuel-air mixing, and finite-rate chemistry effects. This case employs a reduced, 30-species, 231-reaction mechanism proposed by Smith et al. (2020) and presents significant computational challenges due to the three-dimensional nature of the flow, the presence of multiple shocks, and the strong coupling between turbulent mixing and chemical kinetics in the high-speed regime.
@@ 4.4 Choice of Parameters
*More details of this case can be found in Sharma et al. (2024c)
@@ Appendix A Verification of the Kinetics Solver
Figure 7: Accuracy verification of the kinetics solver. (A,B,C) Comparison of ignition delay times for hydrogen-air mixtures at stoichiometric conditions and 1 atm pressure between Cantera reference solutions (solid circles) and our GPU solver using three chemical mechanisms: (A) the 14-species Mevel mechanism (Mével et al., 2009), (B) the reduced 30-species FFCMy mechanism (Smith et al., 2020), and (C) the 53-species GRI-Mech 3.0 (Smith et al., ) mechanism. (D,E,F) Species molar production rate correlations between GPU solver calculations (y axis) and Cantera calculations (x axis) for the same chemical mechanisms. The dashed line represents perfect agreement. Different colors represent different chemical species. $R^{2}$ values are also reported.
@@ Appendix D Performance Variation with Active Cell Count
Figure 10: Single GPU runtime vs active cell count in one grid for the 14-species mechanism (Mével et al., 2009) and the 30-species mechanism (Smith et al., 2020) for the isolated chemistry routine. The solid green vertical line indicates the point at which GPU saturation is reached.

### v2

@@ Abstract
High-speed chemically active flows present significant computational challenges due to their disparate space and time scales, where stiff chemistry often dominates simulation time. While modern supercomputing scientific codes achieve exascale performance by leveraging graphics processing units (GPUs), existing GPU-based compressible combustion solvers face critical limitations in memory management, load balancing, and handling the highly localized nature of chemical reactions. To this end, we present a high-performance compressible reacting flow solver built on the AMReX framework and optimized for multi-GPU settings. Our approach addresses three GPU performance bottlenecks: memory access patterns through column-major storage optimization, computational workload variability via a bulk-sparse integration strategy for chemical kinetics, and multi-GPU load distribution for adaptive mesh refinement applications. The solver adapts existing matrix-based chemical kinetics formulations to multigrid contexts. Using representative combustion applications including hydrogen-air detonations and jet in supersonic crossflow configurations, we demonstrate $2-5\times$ performance improvements over initial GPU implementations with near-ideal weak scaling across 1961-96 NVIDIA H100 GPUs. Roofline analysis reveals substantial improvements in arithmetic intensity for both convection ($\sim 10\times$) and chemistry ($\sim 4\times$) routines, confirming efficient utilization of GPU memory bandwidth and computational resources.
@@
[affil1] organization=Department of Aerospace Engineering, University of Michigan, city=Ann Arbor, state=Michigan 48105, country=USA
@@ 1 Introduction
Other studies have explored GPU usage or adaptive mesh refinement (AMR) techniques for accelerating CFD codes (Sharma et al., 2024a; Bernardini et al., 2021; Lai et al., 2019; Barwey and Raman, 2021; Bielawski et al., 2023; Romero et al., 2020; Witherden et al., 2015). AMR is an approach that helps address the disparate space and time scales of compressible reacting flow simulations on large, yet fine, meshes by resolving only spatio-temporal regions that need higher resolution. For instance, the AMReX framework (Zhang et al., 2019) — a structured-grid AMR coding framework — provides a hardware-agnostic solution for CFD code development, easing code portability to a diverse set of exascale architectures. AMReX-based solvers such as PeleC (Henry de Frahan et al., 2023) have demonstrated remarkable scalability. However, achieving optimal GPU performance demands algorithmic restructuring and hardware-specific tuning beyond CPU-GPU porting strategies provided by AMReX.
@@ 2.2 Graphics Processing Units and Roofline Models
Graphics processing units (GPUs) can significantly accelerate CFD calculations because many underlying algorithms can be cast into a single-instruction, multiple-data (SIMD) pattern, which GPUs are designed to exploit (Flynn, 1972). While CPUs excel at complex control flow and low-latency operations, GPUs contain hundreds of processing cores termed streaming multiprocessors (SMs) designed to execute identical operations on different data elements simultaneously. This hardware design makes them well suited for grid- and particle-based simulations, where the same numerical operations are applied independently across millions of computational elements with minimal interdependence. However, effectively utilizing GPUs requires addressing fundamental differences from CPU programming. A kernel is a function that runs many times in parallel on a GPU. The kernel executes work across many parallel workers, termed threads. The threads work in unison within groups called warps (32 threads per warp on NVIDIA hardware) that must execute identical instructions in lockstep. When threads within a warp take different execution paths – a condition known as thread divergence – performance degrades significantly. Additionally, GPUs achieve peak performance only when memory accesses follow coalesced patterns, where consecutive threads access consecutive memory locations (NVIDIA Corporation, 2025a).
In modern GPUs, performance is predominantly limited by memory bandwidth rather than arithmetic throughput (Nickolls and Dally, 2010). For instance, the NVIDIA H100 GPU (NVIDIA Corporation, 2025d) used in this study achieves peak double-precision performance of 34 TFLOPs/s but has global memory bandwidth of only 3.35 TB/s. This imbalance means that most computational kernels become memory-bound, where the costs of data movement dominate the costs of arithmetic operations. Common optimization strategies include minimizing CPU-GPU data transfers, utilizing on-chip shared memory for data reuse, ensuring coalesced memory access patterns, and maximizing arithmetic intensity—the ratio of floating-point operations to bytes transferred (Schroeder, 2011).
@@
Modern roofline analysis tools, such as NVIDIA Nsight Compute (NVIDIA Corporation, 2025e), automate the collection of performance metrics necessary for roofline analysis. These tools measure kernel execution time, floating-point operations, and memory transactions to calculate both arithmetic intensity and sustained performance. The resulting roofline charts provide immediate visual feedback on performance bottlenecks and guide optimization efforts by clearly indicating whether kernel performance is limited by computational throughput, or other architectural constraints.
@@ 2.3 Overview of AMReX
This section provides a brief overview of the AMReX framework, focusing on the key concepts and features most relevant to our GPU-accelerated combustion solver. For comprehensive documentation of AMReX’s full capabilities, readers are referred to Zhang et al. (2019) and the official documentation (AMReX Development Team, 2025).
@@
The AMReX framework organizes computational cells into rectangular non-overlapping grids at different AMR levels. The coarsest level – level 0 or the base level – contains the base coarse grids that span the entire physical domain, while higher levels contain dynamically generated finer grids that cover only regions meeting user-defined refinement criteria. To ensure efficient multigrid performance and optimal load balancing, AMReX enforces a blocking factor constraint, requiring all grid dimensions to be divisible by a specified power-of-two value (e.g., $2^{k}$ where $k\geq 0$). This constraint enables efficient grid refinement algorithms and facilitates regular memory access patterns, which are essential for achieving optimal GPU performance. From a programming standpoint, a Fortran Array Box (FAB) object efficiently stores the data and attributes for each grid, and FABs at a given AMR level are collectively stored in a multiFAB object (AMReX Development Team, 2025). multiFAB provides a high-level interface for computational scientists to perform efficient parallel operations, load balancing, and communication. Unlike unstructured adaptive methods, this block-structured approach preserves regular data access patterns, which are essential for achieving optimal GPU performance.
@@
Load balancing across multiple GPUs is a key challenge in multigrid solvers. In CPU settings, users can calculate computational costs for each MultiFab and provide these to AMReX’s load balancing algorithms, which then redistribute the grids across the available MPI ranks to achieve balanced workloads AMReX Development Team (2025). However, this strategy becomes limited on GPUs due to the typical one-rank-per-GPU deployment, which drastically reduces the number of redistribution targets compared to CPU runs that may use dozens of ranks per node. Additionally, GPU performance is nonlinear with respect to the number of grids and cells processed, making cost estimation more complex. Load balancing becomes particularly important in scientific codes where physical features – and thus computational costs – can vary dramatically across the domain.
@@ 3.1 Memory-Based Optimizations
AMReX already provides several GPU-based strategies for memory management and access. For instance, the framework manages memory through pre-allocated memory arenas that reuse contiguous memory chunks to eliminate allocation overhead (AMReX Development Team, 2025). By default, amrex.the_arena_is_managed is set to 0, allocating device-only memory. Setting it to 1 enables managed memory accessible from both the CPU and GPU, facilitating easier development but potentially incurring performance penalties due to page faults. To address the penalties, we disabled managed memory and use explicit memory management via The_Async_Arena() for temporary allocations while leveraging AMReX’s arena system for memory pools. We also determined the maximum problem size for a single H100 based on the total device memory requirements for various variables. The variables include, but are not limited to, thermo-fluid variables, chemistry variables, temporary variables for flux computation, and variables for ghost cells.
@@ 3.2 Chemical Kinetics Solver
To reduce kernel launch costs in both bulk and sparse integration, we launch a single kernel across all cells from all grids. This design choice sacrifices some parallelization strategies over others. To guide this explanation, refer to Figure 2, which showcases two implementations of kinetics on GPUs. In a single-grid solver with contiguous memory layout, we could parallelize across both cells and species ($N_{c}\times N_{s}$ threads), enabling the use of GPU-optimized libraries like cuBLAS (NVIDIA Corporation, 2025c) for matrix operations, as seen in Figure 2. This approach was utilized by Barwey and Raman (2021) to achieve superior performance when working with (i) a single medium to large grid (see C) or (ii) a low grid count in a multi-grid setting (see E). This multi-kernel strategy also eliminates warp divergence by checking whether all cells are complete before continuing the while loop iteration.
@@ 4.1 Setup of the Hardware, Software, and Performance Metrics
We obtain all results reported in Section 5 by running the solver on a cluster with NVIDIA H100 GPUs (NVIDIA Corporation, 2025d). The GPUs are peer-to-peer connected using 18 4th-generation NVLINKs, with a total peak bandwidth of 900 GB/s, supporting simultaneous bidirectional communication. Empirical bandwidth tests using sample CUDA scripts (NVIDIA Corporation, 2025b) measured net unidirectional data transfer rates of $\approx 478$ GB/s between any two GPUs. GPUs are connected to Intel(R) Xeon(R) Platinum 8468 CPUs using PCIe 5th Gen communication technologies, with peak bandwidths of 64 GB/s. The simulations were executed with each CPU process bound to a single GPU, as required by the AMReX GPU implementation. Therefore, the number of CPUs used for a simulation equals the number of GPUs.
We compiled two versions of the executable for a given set of inputs. The first version is compiled with level 3 optimizations using the GNU suite of compilers. The second version is a suboptimal compilation with debugging enabled, allowing for the collection of wall timing for various code regions. We also used the debugging version to collect detailed GPU kernel metrics using NVIDIA Nsight Compute (NVIDIA Corporation, 2025e) to derive quantities for the roofline model. These metrics are detailed in Yang et al. (2021).
@@ 4.3 Description of The Combustion Applications
Jet in Supersonic Crossflow (JISCF): This application simulates a sonic Ethylene jet injected into a high-temperature Mach 2.48 crossflow, representing conditions relevant to high-speed propulsion systems. The interested reader may find details of the case in Sharma et al. (2024b). The configuration features complex physics, including compressible turbulence, shock-turbulence interactions, fuel-air mixing, and finite-rate chemistry effects. This case employs a reduced, 30-species, 231-reaction mechanism proposed by Smith et al. (2016) and presents significant computational challenges due to the three-dimensional nature of the flow, the presence of multiple shocks, and the strong coupling between turbulent mixing and chemical kinetics in the high-speed regime.
@@ 4.4 Choice of Parameters
*More details of this case can be found in Sharma et al. (2024b)
@@ Appendix A Verification of the Kinetics Solver
Figure 7: Accuracy verification of the kinetics solver. (A,B,C) Comparison of ignition delay times for hydrogen-air mixtures at stoichiometric conditions and 1 atm pressure between Cantera reference solutions (solid circles) and our GPU solver using three chemical mechanisms: (A) the 14-species Mevel mechanism (Mével et al., 2009), (B) the reduced 30-species FFCMy mechanism (Smith et al., 2016), and (C) the 53-species GRI-Mech 3.0 (Smith et al., ) mechanism. (D,E,F) Species molar production rate correlations between GPU solver calculations (y axis) and Cantera calculations (x axis) for the same chemical mechanisms. The dashed line represents perfect agreement. Different colors represent different chemical species. $R^{2}$ values are also reported.
@@ Appendix D Performance Variation with Active Cell Count
Figure 10: Single GPU runtime vs active cell count in one grid for the 14-species mechanism (Mével et al., 2009) and the 30-species mechanism (Smith et al., 2016) for the isolated chemistry routine. The solid green vertical line indicates the point at which GPU saturation is reached.